Solve: 5x-8-2x=4x-6
Answer by rapaljer(4671)

5x-8-2x=4x-6

Combine like terms:
3x-8=4x-6

Subtract 3x from each side:
3x-3x-8=4x-3x-6
-8=x-6

Next, add +6 to each side:
-8+6=x-6+6
-2=x

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida
Altamonte Springs Campus